# Unveiling the optimization process of Physics Informed Neural Networks: How accurate and competitive can PINNs be?

### v1

@@ 4 Other physics problems
Regarding the pressure, we only need to specify it at a single spacial point $(x_{1},y_{1},z_{1})$ for all times $t$. This is necessary, because one has the freedom to add a function of time to $p$ and get an equivalent solution. Indeed, if we replace $p\rightarrow p+f(t)$ in equation (51) we obtain the same system and the solution is ambiguous. Results for the loss function and the error estimations can be found in figure 11(d) and table 10.
Figure 11: Loss function vs number of iterations for the four problems considered in section 4. Results are obtained by considering Adam with $\mathcal{J}$, Adam + BFGS with $\mathcal{J}$, Adam + BFGS with $\mathcal{J}_{\text{log}}$, and Adam + SSBroyden with $\mathcal{J}$. 11(a): Non-linear Poisson equation; 11(b): Non-linear Schrödinger equation; 11(c): KdV equation; 11(d): Navier-Stokes (Beltrami flow) equation.
@@ 4.1 Comparison with the literature
Wrapping up this section, in Table 3 we offer a direct one-to-one comparison of our findings (marked as TW –this work–) presented earlier with similar problems encountered in the PINNs literature (specific references provided in the second column). We utilized the code provided in the GitHub repository cited in [42] for the KdV equation, incorporating the analytical 2-soliton solution and defining the same training domain for this study. As for the NLSE, we employed our code (as the L-BFGS optimizer used in [2] is the same as the one considered here) to compute the corresponding PINN approximation errors defined in 3.2.
@@
In the current landscape of PINN research, L-BFGS has become a widely used optimizer due to its generally superior performance compared to the Adam optimizer. However, it has demonstrated some limitations, especially when confronted with highly ill-conditioned problems, as evidenced in some studies [44]. Consequently, it needs a significantly higher number of trainable parameters than the BFGS algorithm to achieve comparable precision. Naturally, training the identical network with BFGS is markedly slower compared to L-BFGS, primarily due to the former’s requirement to store and update the inverse Hessian estimate at each iteration. However, the ability to utilize smaller networks with BFGS enables obtaining results with comparable or even higher accuracy than those obtained with larger networks with the L-BFGS algorithm.
@@
Figure 12 illustrates this effect. We show the evolution of $\mathcal{J}$ as a function of the number of iterations (left panel) or training time (right panel) for the NLSE equation. We compare the training process with the L-BFGS and the BFGS algorithms using our smaller neural network (see table 1), and the L-BFGS algorithm with the same NN considered in [2].
@@
As shown by the figure, for identical networks, the L-BFGS algorithm operates approximately four times faster, yet its convergence rate is notably slower. Thus, we note that BFGS proves to be more efficient than L-BFGS for the same number of trainable parameters, both in terms of iterations and training duration. While employing a larger network with L-BFGS enhances accuracy, it does so at the expense of significantly prolonged training periods.
@@
However, the preceding examples in this section illustrate that training with BFGS in conjunction with the MSE loss was not the optimal strategy either. Figure 13 shows again the evolution of $\mathcal{J}$ as a function of the training time. By introducing modifications to either the BFGS algorithm or the loss function, the performance enhancement becomes even more remarkable. With these adjustments, we observe a reduction of many orders of magnitude in the loss function within the same training duration, compared to L-BFGS. This is directly reflected in the errors achieved in each case, as shown in table 3.
@@
TW BFGS $\mathcal{J}$ 2 30 $7.65\times 10^{-7}$
TW SSBroyden $\mathcal{J}$ 2 40 $4.60\times 10^{-9}$
TW BFGS $\mathcal{J}_{\mathrm{log}}$ $2$ 40 $6.86\times 10^{-9}$
NLS $\left(\epsilon_{\mathrm{NN}}^{(u)},\epsilon_{\mathrm{NN}}^{(v)}\right)$
@@
TW BFGS $\mathcal{J}$ 2 40 $(4.32,4.16)\times 10^{-4}$
TW SSBroyden $\mathcal{J}$ 2 40 $(2.42,2.40)\times 10^{-5}$
TW BFGS $\mathcal{J}_{\mathrm{log}}$ $2$ 40 $(3.85,3.38)\times 10^{-5}$
@@
TW BFGS $\mathcal{J}$ 3 30 $4.16\times 10^{-4}$
TW SSBroyden $\mathcal{J}$ 3 30 $6.74\times 10^{-6}$
TW BFGS $\mathcal{J}_{\mathrm{log}}$ $3$ 30 $1.22\times 10^{-5}$
@@
[45] L-BFGS $\mathcal{J}$ 7 50 $(2.54,2.40,2.60)\times 10^{-4}$
TW BFGS $\mathcal{J}$ 2 40 $(1.30,1.59,1.63)\times 10^{-5}$
TW SSBroyden $\mathcal{J}$ 2 40 $(7.24,7.07,7.78)\times 10^{-7}$
TW BFGS $\mathcal{J}_{\mathrm{log}}$ $2$ 40 $(1.27,1.36,1.33)\times 10^{-6}$
Table 3: Comparison between models in the literature (references in the second column) and our models (this work TW) trained with our optimizer and loss function modifications. If an analytical solution exists, we compute the error as the $L_{2}$ relative norm, denoted for a given solution $u$ by $E_{u}$. If not (as in the non-linear Schrödinger equation), we compute the PINN approximation error $\epsilon_{\mathrm{NN}}$ defined in (3.2).
Figure 12: Loss function obtained for the NLSE equation considering the L-BFGS and BFGS algorithms with the network employed in this work (2L40N) for this particular problem (see table 1) and the L-BFGS algorithm for the model suggested in [2] (4L100N), as a function of the number of iterations (12(a)) and the training time (12(b)).
Figure 13: Loss function vs. training time for the NLSE equation obtained with our training process modifications. The ones presented in figure 12(b) are also shown for reference.
@@ Acknowledgments
We acknowledge the support through the grant PID2021-127495NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by the European Union, the Astrophysics and High Energy Physics programme of the Generalitat Valenciana ASFAE/2022/026 funded by MCIN and the European Union NextGenerationEU (PRTR-C17.I1) and the Prometeo excellence programme grant CIPROM/2022/13. JFU is supported by the predoctoral fellowship ACIF 2023, cofunded by Generalitat Valenciana and the European Union through the European Social Fund.
Appendix A Efficient computation of the scaling parameter $\tau_{k}^{(1)}$
@@
$\tau_{k}^{(1)}=\min\left\{1,\frac{-\bm{y}_{k}\cdot\bm{s}_{k}}{\alpha_{k}\bm{s}_{k}\cdot\nabla\mathcal{J}\left(\bm{\Theta}_{k}\right)}\right\},$ (63)
Note that the original suggestion of [18] for the scaling parameter was simply $\tau_{k}^{(1)}=\frac{-\bm{y}_{k}\cdot\bm{s}_{k}}{\alpha_{k}\bm{s}_{k}\cdot\nabla\mathcal{J}\left(\bm{\Theta}_{k}\right)}$, which was indeed motivated by the need to reduce the condition number of (24). However, this choice has shown to be inferior in terms of performance compared to the standard BFGS, when combined with an inexact line search computation of $\alpha_{k}$, as shown in [46] and was confirmed by our own analysis. Instead, we followed the suggestion introduced in [17], which is a simple modification of the original one, but ensures theoretically super-linear convergence with inexact line searches. In fact, it can be thought as a switch between the standard BFGS algorithm ($\tau_{k}=1$) and the self-scaled BFGS algorithm of [18]. Curiously, we found than in the majority of the training iterations, the value of $\tau_{k}=1$ is selected, with only a small portion of them benefitting from the scaling. Still, this was sufficient to produce considerable improvements in the optimization process.
Appendix B Self-scaled Broyden optimization algorithm
The intermediate variables $\theta_{k}$, $\sigma_{k}$ and $a_{k}$ defined in [15] and used in equations (20), (21) are given by
$h_{k}$ $=\frac{\bm{y}_{k}\cdot H_{k}\bm{y}_{k}}{\bm{y}_{k}\cdot\bm{s}_{k}},$ (65)
$a_{k}$ $=h_{k}b_{k}-1$ (66)
$c_{k}$ $=\sqrt{\frac{a_{k}}{a_{k}+1}},$ (67)
$\rho_{k}^{-}$ $=\min\left(1,h_{k}\left(1-c_{k}\right)\right),$ (68)
$\theta_{k}^{-}$ $=\frac{\rho_{k}^{-}-1}{a_{k}},$ (69)
$\theta_{k}^{+}$ $=\frac{1}{\rho_{k}^{-}},$ (70)
$\theta_{k}$ $=\max\left(\theta_{k}^{-},\min\left(\theta_{k}^{+},\frac{b_{k}}{1-b_{k}}\right)\right)$ (71)
$\sigma_{k}$ $=1+a_{k}\theta_{k}.$ (72)
@@ Appendix C Error analysis
We present tables with error measurements for the various problems that we solve.
For each problem, we repeat the training process for each modification of the optimization algorithm or the loss function using different random initialization for the trainable parameters and the training set. The experiments are labeled with a particular number, indicating a given seed for the random number generator. We calculate the final value of the MSE loss and the errors measured in a given set of tests points. If there exists an analytical solution to the problem, say $u_{\mathrm{an}}$, then we define the error of order $p$ at the PINN approximation $u$ as

### v2

@@ 5 Other physics problems
Figure 20: Convergence plots for the 3D Navier-Stokes equation. Left panel: evolution of the loss function. Right panel: evolution of the relative $L_{2}$ error with respect to the reference solution on a fixed grid.
Figure 21: Solution comparison for the 3D Navier-Stokes equation. Left panels: reference solution. Middle panels: PINN prediction obtained with the SSBroyden algorithm and the standard MSE loss. Right panels: absolute difference between the two. The plots show cuts of the $x-y$ plane at $z=0.5$ and $t=1$. The first three rows show the velocity components. The fourth row shows the pressure.
Regarding the pressure, we only need to specify it at a single spacial point $(x_{1},y_{1},z_{1})$ for all times $t$. This is necessary because one has the freedom to add a function of time to $p$ and get an equivalent solution. Indeed, if we replace $p\rightarrow p+f(t)$ in equation (66) we obtain the same system and the solution is ambiguous.
Figure 20 and table 5 show the results for the loss function and the error norms. Figure 21 shows colormaps in a given cut of the components of the velocity and the pressure predicted with the PINN using the self-scaled Broyden algorithm, together with the analytical ones and the absolute difference between them.
Lid-driven cavity (LDC)
To conclude, we present the results obtained for the 2D lid-driven cavity flow treated in recent studies [24, 76]. This problem is governed by the stationary momentum equation
$\left(\bm{u}\cdot\nabla\right)\bm{u}=-\nabla p+\nu\nabla^{2}\bm{u},$ (77)
together with the continuity equation for an incompressible flow
$\nabla\cdot\bm{u}=0,$ (78)
where $\bm{u}=\left(u(x,y),v(x,y)\right)$ is the vector velocity and $p$ denotes the pressure. The coefficient $\nu$ is commonly expressed in terms of the Reynolds number $\mathrm{Re}$ by setting $\nu=1/\mathrm{Re}$.
The domain is the square $\left[0,1\right]^{2}$. The boundary conditions for this problem are
$\bm{u}_{b}(x,y)=\begin{cases}&\left(u_{T}(x),0\right),\quad y=1,\\ &\left(0,0\right),\quad\mathrm{otherwise}\end{cases}$ (79)
where $u_{T}(x)$ must equal $1$ across most of the top edge of the domain and decay smoothly to zero at the corners to meet the Dirichlet boundary conditions in the horizontal direction. More concretely, we consider the same function as in [24]
$u_{T}(x)=1-\frac{\cosh{\left(C_{0}\left(x-0.5\right)\right)}}{\cosh{\left(0.5C_{0}\right)}}$ (80)
where $C_{0}=50$. Following [76, 82], we employ the stream function formalism, solving for $\psi$ instead of the two components of the velocity. The pressure is the other output of the network and is set to zero at a particular point in the domain to avoid ambiguities in the solution. The total loss function is then given in this case by
$\mathcal{J}=\mathcal{J}_{\mathrm{PDE}}+\frac{\lambda}{N_{b}}\lVert\bm{u}-\bm{u}_{b}\rVert_{(x,y)\in\mathcal{B}}^{2},$ (81)
where $\mathcal{B}$ is the boundary of $\left[0,1\right]^{2}$, and $\mathcal{J}_{\mathrm{PDE}}$ is the sum of the residuals corresponding to the two momentum equations, constructed with (2) by identifying $\mathcal{L}_{u}=\frac{\partial}{\partial t}-\nu\nabla^{2}u$, $G_{u}=-\left(\bm{u}\cdot\nabla\right)u-\frac{\partial p}{\partial x}$ and $\mathcal{L}_{v}=\frac{\partial}{\partial t}-\nu\nabla^{2}v$, $G_{v}=-\left(\bm{u}\cdot\nabla\right)v-\frac{\partial p}{\partial y}$. The following results are obtained by setting $\mathrm{Re}=1000$. The parameter $\lambda$ that controls the contribution of the boundary conditions is set to $10$.
Results for the loss function are shown in figure 22. Figure 23 shows the solution for the components and the modulus of the velocity aquired with the self-scaled Broyden algorithm. Since no analytical or high-precision numerical solution is available, we estimate the error according to equation (91) as $\epsilon_{\mathrm{NN}}\approx 1.3\times 10^{-4}$, obtained with the self-scaled Broyden algorithm. The standard BFGS algorithm on the other hand achieves, if trained against the standard MSE loss $\mathcal{J},$ a value of $\epsilon_{\mathrm{NN}}\approx 1.0\times 10^{-3}$, whereas if the logarithm of the MSE loss $\mathcal{J}_{\mathrm{log}}$ is employed, this value is reduced up to $\epsilon_{\mathrm{NN}}\approx 1.6\times 10^{-4}$.
Figure 22: Loss functions for the lid-driven cavity problem.
Figure 23: PINN predictions for the lid-driven cavity problem. The plots show the components of the vector velocity $\bm{u}$ and its modulus. The lines correspond to contours of the quantities represented in each plot.
@@ 5.1 Comparison with the literature
Wrapping up this section, in table 5 we offer a direct one-to-one comparison of our findings (marked as TW –this work–) presented earlier with similar problems encountered in the PINNs literature (specific references provided in the second column). We utilized the code provided in the GitHub repository cited in [100] for the KdV equation, incorporating the analytical 2-soliton solution and defining the same training domain for this study. As for the NLS equation, we employed our code (as the L-BFGS optimizer used in [1] is the same as the one considered here) to compute the corresponding PINN approximation errors defined in 4.2. Finally, for the Allen-Cahn equation we have employed the code given in the Github repository cited in [102].
@@
In the current landscape of PINN research, L-BFGS has become a widely used optimizer due to its generally superior performance compared to the Adam optimizer. However, it has demonstrated some limitations, especially when confronted with highly ill-conditioned problems, as evidenced in some studies [105]. Consequently, it needs a significantly higher number of trainable parameters than the BFGS algorithm to achieve comparable precision. Although training two identical networks with BFGS would be markedly slower compared to L-BFGS, primarily due to the former’s requirement to store and update the inverse Hessian estimate at each iteration, with BFGS one can use much smaller networks with higher accuracy.
@@
Figure 24 illustrates this effect. We show the evolution of $\mathcal{J}$ as a function of the number of iterations (left panel) or training time (right panel) for the NLS equation. We compare the training process with the L-BFGS and the BFGS algorithms using our smaller neural network (see table 1), and the L-BFGS algorithm with the same NN considered in [1].
@@
For identical networks, L-BFGS runs about four times faster, but its convergence rate is significantly slower. Using a larger network with L-BFGS improves accuracy but significantly increases training time. Overall, to achieve a given accuracy, BFGS proves more efficient than L-BFGS for the same number of parameters, both in terms of iterations and total training time.
@@
However, the preceding examples in this section illustrate that training with BFGS in conjunction with the MSE loss was not the optimal strategy either. By introducing modifications to either the BFGS algorithm or the loss function, the performance enhancement becomes even more remarkable, as shown in Figure 25.
With these adjustments, we observe a reduction of several orders of magnitude in the loss function within the same training duration, compared to L-BFGS. This is directly reflected in the errors of the solution achieved in each case, summarized in table 5. The errors computed in this work are averaged over multiple training runs, employing different random initializations of the trainable parameters to ensure robustness. Our results are presented in the format (mean ± standard deviation), to illustrate both, the average error and the variability across the different trials. In all cases, our solutions are 1-2 orders of magnitude more precise than the references in the literature, although we are using significantly smaller networks (see columns 5-6 in the table).
Table 5: Comparison between models in the literature (references in the second column) and our models in this work (TW), with modifications in the optimizer and loss function. Errors for a given variable $x$ are denoted by $E_{x}$ and are computed as the $L_{2}$ relative norm of the difference with a reference analytical or numerical solution, (see equation (89)).
@@
2DH $(1,4)$ $E_{u}$
[57] L-BFGS $\mathcal{J}$ 6 128 $8.21\times 10^{-6}$
TW SSBroyden $\mathcal{J}$ 2 20 $(6\pm 4)\times 10^{-8}$
TW BFGS $\mathcal{J}_{\mathrm{log}}$ $2$ 20 $(3.6\pm 0.2)\times 10^{-7}$
@@
TW SSBroyden $\mathcal{J}$ 2 30 $(3\pm 1)\times 10^{-9}$
TW BFGS $\mathcal{J}_{\mathrm{log}}$ $2$ 30 $(6\pm 3)\times 10^{-9}$
NLS $\left(E_{u},E_{v}\right)$
@@
TW SSBroyden $\mathcal{J}$ 2 40 $(5\pm 1,8\pm 2)\times 10^{-6}$
TW BFGS $\mathcal{J}_{\mathrm{log}}$ $2$ 40 $(1\pm 0.1,1.5\pm 0.3)\times 10^{-5}$
@@
TW SSBroyden $\mathcal{J}$ 3 30 $(6\pm 0.7)\times 10^{-6}$
TW BFGS $\mathcal{J}_{\mathrm{log}}$ $3$ 30 $(1.2\pm 0.4)\times 10^{-5}$
1DB $E_{u}$
[102] Adam $\mathcal{J}$ 6 128 $4.8\times 10^{-4}$
TW SSBroyden $\mathcal{J}$ 3 20 $(2.9\pm 0.4)\times 10^{-6}$
TW BFGS $\mathcal{J}_{\mathrm{log}}$ $3$ 20 $(5\pm 2)\times 10^{-6}$
AC $E_{u}$
[102] Adam $\mathcal{J}$ 6 128 $1.45\times 10^{-5}$
TW SSBroyden $\mathcal{J}$ 3 30 $(2.2\pm 0.7)\times 10^{-6}$
TW BFGS $\mathcal{J}_{\mathrm{log}}$ $3$ 30 $(9.7\pm 0.8)\times 10^{-6}$
@@
[106] L-BFGS $\mathcal{J}$ 7 50 $(2.54,2.40,2.60)\times 10^{-5}$
[107] L-BFGS $\mathcal{J}$ 5 50 $(2.64,4.35,2.74)\times 10^{-5}$
TW SSBroyden $\mathcal{J}$ 2 40 $(7.3\pm 0.9,7.1\pm 0.4,7.8\pm 0.1)\times 10^{-7}$
TW BFGS $\mathcal{J}_{\mathrm{log}}$ $2$ 40 $(1.3\pm 0.1,1.4\pm 0.1,1.3\pm 0.1)\times 10^{-6}$
Figure 24: Loss function obtained for the non-linear Schrödinger equation considering the L-BFGS and BFGS algorithms with the network employed in this work (2L40N) and the L-BFGS algorithm for the model suggested in [1] (4L100N), as a function of the number of iterations (24(a)) and the training time (24(b)).
Figure 25: Loss function vs. training time for the non-linear Schrödinger equation obtained with our training process modifications. The results presented in figure 24(b) are also shown for reference.
@@ Acknowledgments
We acknowledge the support through the grant PID2021-127495NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by the European Union, the Astrophysics and High Energy Physics programme of the Generalitat Valenciana ASFAE/2022/026 funded by MCIN and the European Union NextGenerationEU (PRTR-C17.I1), and the Prometeo excellence programme grant CIPROM/2022/13. JFU is supported by the predoctoral fellowship ACIF 2023, cofunded by Generalitat Valenciana and the European Union through the European Social Fund.
Appendix A Derivation of the Grad-Shafranov equation
In axisymmetry, the magnetic field $\bm{B}$ can be described in terms of two poloidal and toroidal stream functions $\mathcal{P}$ and $\mathcal{T}$ as
where $\bm{e}_{\phi}$ is the unit vector in the $\phi$ direction.
$\nabla\mathcal{P}\times\nabla\mathcal{T}=0,$ (83)
which implies that $\mathcal{T}=\mathcal{T}(\mathcal{P})$, and equation (28). The magnetic field components can be recovered from $\mathcal{P}$ and $\mathcal{T}$ through the following relations:
$B_{r}$ $=-q^{2}\frac{\partial\mathcal{P}}{\partial\mu},$ (84)
$B_{\theta}$ $=\frac{q^{3}}{\sqrt{1-\mu^{2}}}\frac{\partial\mathcal{P}}{\partial q},$ (85)
$B_{\phi}$ $=\frac{q}{\sqrt{1-\mu^{2}}}\mathcal{T}$ (86)
Appendix B Efficient computation of the scaling parameter $\tau_{k}^{(1)}$
@@
$\tau_{k}^{(1)}=\min\left\{1,\frac{-\bm{y}_{k}\cdot\bm{s}_{k}}{\alpha_{k}\bm{s}_{k}\cdot\nabla\mathcal{J}\left(\bm{\Theta}_{k}\right)}\right\}.$ (88)
Note that the original suggestion of [74] for the scaling parameter was $\tau_{k}^{(1)}=\frac{-\bm{y}_{k}\cdot\bm{s}_{k}}{\alpha_{k}\bm{s}_{k}\cdot\nabla\mathcal{J}\left(\bm{\Theta}_{k}\right)}$, which was indeed motivated by the need to reduce the condition number of (35). However, this choice has shown to be inferior in terms of performance compared to the standard BFGS, when combined with an inexact line search computation of $\alpha_{k}$, as shown in [108] and was confirmed by our own analysis. Instead, we followed the suggestion introduced in [69], which is a simple modification of the original one, but ensures theoretically super-linear convergence with inexact line searches. It can be thought of as a switch between the standard BFGS algorithm ($\tau_{k}=1$) and the self-scaled BFGS algorithm of [74]. Curiously, we found that in the majority of the training iterations, the value of $\tau_{k}=1$ is selected, with only a small portion of them benefitting from the scaling. Still, this was sufficient to produce considerable improvements in the optimization process.
@@ Appendix C Error analysis
If analytical or high-fidelity numerical solutions to the problem, denoted as $u_{\mathrm{an}}$, exist, the relative error of the PINN approximation $u$ is defined as:
$E_{u}^{(2)}=\frac{\lVert u-u_{\mathrm{an}}\rVert_{2}}{\lVert u_{\mathrm{an}}\rVert_{2}},$ (89)
where .$\lVert.\rVert_{2}$ indicates the $L_{2}$-norm evaluated in a given set of test points. When these are not available, we follow the method proposed in [72]. To demonstrate this procedure, we use the 2D non-linear GS equation as an example. We begin by generating a uniform test grid $\left\{q_{i},\mu_{j}\right\}_{i,j=1}^{N_{0}}$, and then evaluate the PINN solution $\mathcal{P}_{ij}=\mathcal{P}(q_{i},\mu_{j},\bm{\Theta})$ at all points on the grid through a forward pass. Then we calculate the derivatives of $\mathcal{P}_{ij}$ using a second-order finite differences scheme and construct a finite difference version of the PDE (28). We define the normalized $L_{2}$-error of the discretized PDE as